OpenStudy (anonymous):

The equation of line AB is y = 2x + 4. Write an equation of a line parallel to line AB in slope-intercept form that contains point (3, -2). (4 points) Question 4 options: 1) y = -2x + 4 2) y = x - 3) y = x - 4) y = 2x - 8 *Help please

OpenStudy (jdoe0001):

hmm well.. what's the slope of y = 2x + 4 anyway?

OpenStudy (anonymous):

1) y = -2x + 4 2) y = -1/2 x - 1/2 3) y = 1/2 x - 7/2 4) y = 2x - 8

OpenStudy (jdoe0001):

ok... so... what did you get for the slope of y = 2x + 4 again?

OpenStudy (anonymous):

is it D ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

Point slope form is y-y1= m(x-x1). Parallel lines have the same slopes. Plug in 2 for m and plug in the coordinates. y-(-2) = 2(x-3). y+2 = 2x-6. y = 2x-8.

OpenStudy (anonymous):

so it's D? @ridoleor

OpenStudy (anonymous):

Yes it is.

OpenStudy (anonymous):

thank you @ridoleor